4.) Suppose that that there are only two people in society, Mark and Judy, who must split a fixed amount of income of $300. Mark’s Utility function is UM and income is IM. Judy’s Utility function is UJ and income is IJ. Suppose that

Let the Social Welfare function be

W = UM + UJ

What distribution of the total income between Mark and Judy maximizes social welfare?
Answer #1

We have Im + Ij = 100

To maximize welfare we equate the marginal utility of Mark and Judy from Income.

MUm = dUm/dI = 50Im-0.5

MUj = dUj/dI = 100Ij-0.5

Now, MUm = MUj

50Im-0.5 = 100Ij-0.5

Im-0.5 = 2Ij-0.5

2Im0.5 = Ij0.5

4Im = Ij

We have Im+Ij =100, Ij = 100-Im

4Im = 100-Im

Im = 20

Hence Ij = 80